# Active Human Feedback Collection via Neural Contextual Dueling Bandits

### v1

@@ B.2 Additional Results
Figure 2: Performance comparison under different input dimensionalities, i.e., 20-dimensional input (first row), 40-dimensional input (second row). The arm size is 10. We are using Square function for these results.
@@
Figure 3: Performance comparison under different numbers of arms. The first row is when the arm size is 5, and the second row is when the arm size is 10, the third row is when the arm size is 15. The dimensionality of input is 20. We are using Square function for these results.
[…]
(a) Suboptimality Gap
(b) MAE
(c) Average Regret
(d) Suboptimality Gap
(e) MAE
(f) Average Regret
Figure 4: Performance of Neural-ADB (UCB) under different input dimensions (first row) and numbers of arms (second row). We are using Square function for these results.
@@
Figure 5: Performance of Neural-ADB (TS) under different input dimensions (first row) and numbers of arms (second row). We are using Square function for these results.

### v2

@@ B.2 Additional Experimental Results
Figure 4: Performance comparison across different input dimensions $d$: $d=20$ (first row) and $d=40$ (second row). We set the number of arms to $10$ and use the Square function for all experiments.
@@
Figure 5: Performance comparison across different numbers of arms $K$: $K=5$ (top row), $K=10$ (middle row), and $K=15$ (bottom row). We set the input dimension to $20$ and use the Square function for all experiments.
Performance vs. neural network size.   To investigate how performance varies with different neural network (NN) sizes, we used the Square and Cosine functions defined in the paper. We varied either the number of layers (with width = 32) or the width of the NN (with 2 layers), while keeping all other variables consistent with those in the paper. As shown in Fig. 6, we observed that selecting the appropriate size of NN is crucial for the given problem. Using a large NN for a simple problem leads to poor performance due to high bias in the estimation, while a smaller NN may not accurately be able to estimate the complex non-linear function.
@@
Figure 6: We compare performance across different neural network widths (first row) and numbers of hidden layers (second row), using the Square function in all experiments. All other parameters are kept fixed, except that the width is set to $32$ when varying the number of layers.
B.3 Computational Efficiency.
To discuss the computational efficiency of Neural-ADB, we follow the approach of (Verma et al., 2025) and consider the following two key aspects: size of the neural network and then the number of contexts and arms.
Size of the neural network.   The primary computational cost in Neural-ADB arises from the neural network (NN) used to approximate the latent non-linear reward function. Given a context-arm feature vector of dimension $d$, an NN with $D$ hidden layers and $w$ neurons per layer incurs an inference cost of $O(dw+Dw^{2}+w)$ per context-arm pair. The total number of parameters in the NN is $p=dw+Dw^{2}+w$, and the training time per iteration is $O(\mathcal{E}\mathcal{P}Dw^{2})$, where $\mathcal{E}$ is the number of training epochs and $\mathcal{P}$ is the number of observed context-arm pairs. Choosing an appropriate NN size is critical, as NNs that are too small may fail to accurately approximate the underlying non-linear reward function, while excessively large NNs can result in substantial training and inference overhead.
Number of contexts and arms.   Let $K$ denote the number of arms and $p$ the total number of NN parameters. Since Neural-ADB uses NN gradients as context-arm features, the cost of computing gradients for all arms per context is $O(K^{2}dp)$, where $d$ is the dimension of the context-arm feature vector. The cost of computing reward estimates and confidence terms for all context-arm pairs is $O(K^{2}p)$ and $O(K^{2}p^{2})$, respectively. For arm selection, the first selection step requires $O(Kp+K)$, consisting of reward estimation for all arms ($O(Kp)$) and then identifying the arm with the highest estimated reward ($O(K)$). The second arm selection incurs a cost of $O(Kp+(K-1)p^{2})$, including reward estimation $O(Kp)$ and confidence term computation $O((K-1)p^{2})$ relative to the first selected arm. Thus, the total computational cost for selecting a pair of arms per context is $O(K^{2}dp+K^{2}p^{2})$. Since each context-arm pair is independent, gradients, reward estimates, and optimistic terms can be computed in parallel, reducing the overall cost to $O(dp+p^{2})$ for each iteration.